# Occupied Processes: Going with the Flow

### v4

@@ Abstract
In the financial applications, we demonstrate that occupation flows provide unified Markovian lifts for exotic options and variance instruments, allowing financial institutions to price and manage derivatives books with a single numerical solver. We then explore avenues in financial modeling where volatility is driven by the occupied process. In particular, we propose the local occupied volatility (LOV) model which not only calibrates to European vanilla options but also offers the flexibility to capture stylized facts of volatility and fit other instruments. We also present an extension of forward variance models that leverages the entire forward occupation surface.
@@ 1 Introduction
This work originates from an identity in Revuz and Yor [74, Chapter VI],
@@
where $X$ is Brownian motion and $L^{x}$ its local time process at $x$. We refer to $L_{t}^{X_{t}}$ as the spot local time. To prove (1), let $x^{*}$ be a maximizer [^2] The maximum is achieved since $x\mapsto L_{T}^{x}$ has compact support and admits a continuous version [74, Chapter VI]. of $x\mapsto L^{x}_{T}$ and define the random time $\tau^{*}=\sup\{t\leq T\ |\ X_{t}=x^{*}\}$. Then $\sup_{x}L^{x}_{T}=L_{\tau^{*}}^{X_{\tau^{*}}}$, which shows that $\sup_{x}L^{x}_{T}\leq\sup_{t\leq T}L_{t}^{X_{t}}$. The reversed inequality is immediate. One question arises: what if the anticipative strategy $\tau^{*}$ is replaced by stopping times? I.e., what is the optimal rule and value of the optimal stopping (OS) problem,
@@
$\operatorname{\mathcal{O}}_{t}(A)=\int_{0}^{t}\mathds{1}_{A}(X_{s})d\langle X\rangle_{s},\quad A=\text{Borel set of }\operatorname{\mathbb{R}}.$ (3)
See [50], [74, Chapter VI], and Section 2. The occupation flow $\operatorname{\mathcal{O}}=(\operatorname{\mathcal{O}}_{t})_{t\geq 0}$ is a measure-valued process adapted to the filtration of $X$. Moreover, the occupied process $(\operatorname{\mathcal{O}},X)$ is Markov provided that $X$ is Markov.
@@ 1.1 Contributions
$dX_{t}$ $=b(\operatorname{\mathcal{O}}_{t},X_{t})dt+\sigma(\operatorname{\mathcal{O}}_{t},X_{t})dW_{t},\qquad d\operatorname{\mathcal{O}}_{t}=\delta_{X_{t}}\sigma(\operatorname{\mathcal{O}}_{t},X_{t})^{2}dt$ (6)
$dX_{t}$ $=b(\tilde{\operatorname{\mathcal{O}}}_{t},X_{t})dt+\sigma(\tilde{\operatorname{\mathcal{O}}}_{t},X_{t})dW_{t},\qquad\ d\tilde{\operatorname{\mathcal{O}}}_{t}=\delta_{X_{t}}dt$ (7)
@@
$\Big{(}\partial_{\mathfrak{o}}+b\partial_{x}+\frac{\sigma^{2}}{2}\partial_{xx}\Big{)}u(\mathfrak{o},x)=0,\;\mathfrak{o}(\operatorname{\mathbb{R}})<T,\quad u(\mathfrak{o},x)=\varphi(\mathfrak{o},x),\;\mathfrak{o}(\operatorname{\mathbb{R}})=T.$ (8)
@@
Proving the regularity of solutions of path-dependent PDEs is notoriously hard [44, 70] (see also [84, Chapter 11]) and still a subject of active research. In a recent work, Bouchard and Tan [17] study linear parabolic PPDEs for functionals depending on $X_{t}$ and $I_{t}=\int_{0}^{t}X_{s}dA_{s}$ for some continuous function $A$ of finite variation. Note that their setting relates to ours as the integral process $I$ can be retrieved from the occupation measure in some situations. Indeed, if $A_{t}\mathrel{\mathop{\ordinarycolon}}=\langle X\rangle_{t}$ is deterministic (e.g., when $X$ is a Gaussian martingale) and fulfills the conditions in [17], then $I_{t}=\int_{\operatorname{\mathbb{R}}}x\operatorname{\mathcal{O}}_{t}(dx)$.
@@
$\textnormal{v}(t,\omega)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\varphi(\operatorname{\mathcal{O}}_{T},X_{T})|\operatorname{\mathcal{F}}_{t}](\omega),\quad\omega\in\operatorname{\mathcal{D}}([0,T],\operatorname{\mathbb{R}}),$
@@ 1.2 Applications
Sections 4, 5, and 6 are dedicated to financial applications. We shall see that the occupation measure is ever-present, particularly in the eponymous occupation time derivatives [34, 62, 24], timer options, and corridor variance swaps. Other examples include Asian options contingent upon the running average or Parisian options depending on the time spent by the asset above/below a barrier [24, 62]. It is a challenge for financial institutions to consistently price and manage complex derivatives books, and we show in 4 that occupation flows offer a unified framework for this task.
We then explore application of occupied processes in volatility modeling. In Section 5, we introduce the local occupied volatility (LOV) model that lies conveniently between local volatility [38] and path-dependent volatility [52, 53, 56, 59]. This model is designed to both fit the prices of (European) vanilla options and provide additional flexibility to capture stylized facts of volatility and calibrate to other liquid instruments, such as VIX options or American put options in the case of non-dividend-paying stocks. We also propose a generalization of forward variance models in Section 6, opening the doors to nonlinear variance instruments contingent on the floating leg of corridor variance swaps.
@@ 2 Occupied Processes
$\operatorname{\mathcal{O}}_{t}(\omega)\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}})\to\operatorname{\mathbb{R}},\quad\operatorname{\mathcal{O}}_{t}(\omega)(A)=\int_{0}^{t}\mathds{1}_{A}(X_{s})d\langle X\rangle_{s}(\omega),\quad A\in\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}}),\quad\omega\in\Omega.$ (9)
@@
    $\operatorname{\mathcal{O}}_{\tau}(A)=\operatorname{\mathcal{O}}_{\varsigma}(A)+\operatorname{\mathcal{O}}_{\varsigma,\tau}(A),\quad\operatorname{\mathbb{Q}}-a.s.,$ (10)
@@
Evidently, the (strong) time additivity of $\operatorname{\mathcal{O}}$ carries over to $L$, that is $L_{\tau}=L_{\varsigma}+L_{\varsigma,\tau}$ with $L_{\varsigma,\tau}\mathrel{\mathop{\ordinarycolon}}=\frac{d\operatorname{\mathcal{O}}_{\varsigma,\tau}}{d\lambda}$ for all stopping times $\varsigma\leq\tau$. The local time is in fact central in the theory of additive functionals [14], [74, Chapter X]. For local martingales, it is well-known that the process $L_{t}=(L_{t}^{x})_{x\in\operatorname{\mathbb{R}}}$ (indexed by $x$) admits a version that is $\alpha$-Hölder continuous for all $\alpha<1/2$. For general semimartingales, $x\mapsto L_{t}^{x}$ may only shown to be càdlàg [74, Chapter VI]. Finally, note that $\operatorname{\mathcal{O}}_{t}(\operatorname{\mathbb{R}})=\langle X\rangle_{t}$, i.e., the total mass of $\operatorname{\mathcal{O}}$ coincides with the stochastic clock of $X$.
@@ 2.1 Calendar Time Occupation Flow
$\tilde{\operatorname{\mathcal{O}}}_{t}(A)\mathrel{\mathop{\ordinarycolon}}=\int_{0}^{t}\mathds{1}_{A}(X_{s})ds,\quad A\in\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}}),\quad t\in[0,T],$ (14)
@@ 2.3 Occupation Functional and Markov Property
$W_{\theta}\mathrel{\mathop{\ordinarycolon}}=X_{\tau_{\theta}},\qquad\tau_{\theta}=\inf\{t\in[0,T]\ \mathrel{\mathop{\ordinarycolon}}\ \langle X\rangle_{t}>\theta\},\qquad\theta\leq\Theta\mathrel{\mathop{\ordinarycolon}}=\langle X\rangle_{T},$
@@
Note that $\theta_{\tau}$ is a stopping time under the Brownian filtration $\operatorname{\mathbb{F}}^{W}=(\operatorname{\mathcal{F}}_{\tau_{\theta}})_{\theta\leq\Theta}$. This observation will be particularly useful in Section 7 to derive simple formulation of optimal stopping problems. See also [71, Chapter IV].
@@ 2.4 Metrics on Space of Measures
$\operatorname{\mathfrak{m}}_{p}(\mathfrak{o},\mathfrak{o}^{\prime})\mathrel{\mathop{\ordinarycolon}}=\lVert\mathfrak{o}-\mathfrak{o}^{\prime}\rVert_{p},\quad\lVert\mathfrak{o}\rVert_{p}\mathrel{\mathop{\ordinarycolon}}=\left\lVert\frac{d\mathfrak{o}}{d\lambda}\right\rVert_{L^{p}(\operatorname{\mathbb{R}})},\quad\mathfrak{o},\mathfrak{o}^{\prime}\in\operatorname{\mathcal{M}}^{\lambda},\quad p\in[1,\infty].$ (18)
@@
$\operatorname{\mathfrak{m}}_{p}(\operatorname{\mathcal{O}}_{s},\operatorname{\mathcal{O}}_{t})=\lVert\operatorname{\mathcal{O}}_{s,t}\rVert_{p}=\lVert L_{s,t}\rVert_{L^{p}(\operatorname{\mathbb{R}})},\quad s<t.$
@@ 3.1 Occupation Derivative
$df(\operatorname{\mathcal{O}}_{t})=\delta_{\mathfrak{o}}f(\operatorname{\mathcal{O}}_{t})\cdot d\operatorname{\mathcal{O}}_{t},\quad f\in\operatorname{\mathcal{C}}^{1}(\operatorname{\mathcal{M}}),$ (19)
@@
$\phi\cdot\mathfrak{o}\mathrel{\mathop{\ordinarycolon}}=\int_{\operatorname{\mathbb{R}}}\phi(x)\mathfrak{o}(dx),\quad\phi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}},\quad\mathfrak{o}\in\operatorname{\mathcal{M}}.$ (20)
@@
$\delta_{\mathfrak{o}}(\mathfrak{o}\mapsto\phi\cdot\mathfrak{o})\equiv\phi,\quad\forall\phi\in\operatorname{\mathcal{C}}(\operatorname{\mathbb{R}}).$ (22)
@@
$\partial_{\mathfrak{o}}f(\mathfrak{o},x)=\begin{cases}\delta_{\mathfrak{o}}f(\mathfrak{o})(x),\quad&f\in\operatorname{\mathcal{C}}^{1}(\operatorname{\mathcal{M}}),\\[4.25006pt] \delta_{\mathfrak{o}}f(\mathfrak{o},x)(x),\quad&f\in\operatorname{\mathcal{C}}^{1,0}(\operatorname{\mathcal{M}}\times\operatorname{\mathbb{R}}),\end{cases}$ (24)
@@ 3.2 Itô’s Formula
where $\nabla,\nabla^{2}$ denote the gradient and Hessian operator. The occupation derivative is again given by $\partial_{\mathfrak{o}}f(\mathfrak{o},x)=\delta_{\mathfrak{o}}f(\mathfrak{o},x)(x)$. A general Itô’s formula with arbitrary stochastic clock can be found in [77, Proposition 3.3].
@@ 3.2.1 Connection with the Functional Itô Calculus
$d\textnormal{f}(t,\omega)=\Delta_{t}\textnormal{f}(t,\omega)dt+\Delta_{x}\textnormal{f}(t,\omega)dX_{t}(\omega)+\frac{1}{2}\Delta_{xx}\textnormal{f}(t,\omega)d\langle X\rangle_{t}(\omega),\quad\operatorname{\mathbb{Q}}-a.s.,$ (32)
@@ 3.2.2 Connection with the Malliavin Calculus
$D_{t}\textnormal{f}(\omega)\mathrel{\mathop{\ordinarycolon}}=\frac{d}{dh}\textnormal{f}(\omega+h\mathds{1}_{[t,T]})\Big{|}_{h=0}=\int_{t}^{T}\partial_{x\mathfrak{o}}f(\operatorname{\mathcal{O}}_{T}(\omega),\omega_{s})ds,\quad\operatorname{\mathbb{Q}}-a.s.$ (34)

### v5

@@ Abstract
In the financial applications, we demonstrate that occupation flows provide unified Markovian lifts for exotic options and variance instruments, allowing financial institutions to price and manage derivatives books with a single numerical solver. We then explore avenues in financial modeling where volatility is driven by the occupied process. In particular, we propose the local occupied volatility (LOV) model which not only calibrates to European vanilla options but also offers the flexibility to capture stylized facts of volatility or fit other instruments. We also present an extension of forward variance models that leverages the entire forward occupation surface.
@@ 1 Introduction
This work originates from an identity in Revuz and Yor [75, Chapter VI],
@@
where $X$ is Brownian motion and $L^{x}$ its local time process at $x$. We refer to $L_{t}^{X_{t}}$ as the spot local time. To prove (1), let $x^{*}$ be a maximizer [^2] The maximum is achieved since $x\mapsto L_{T}^{x}$ has compact support and admits a continuous version [75, Chapter VI]. of $x\mapsto L^{x}_{T}$ and define the random time $\tau^{*}=\sup\{t\leq T\ |\ X_{t}=x^{*}\}$. Then $\sup_{x}L^{x}_{T}=L_{\tau^{*}}^{X_{\tau^{*}}}$, which shows that $\sup_{x}L^{x}_{T}\leq\sup_{t\leq T}L_{t}^{X_{t}}$. The reversed inequality is immediate. One question arises: what if the anticipative strategy $\tau^{*}$ is replaced by stopping times? I.e., what is the optimal rule and value of the optimal stopping (OS) problem,
@@
$\operatorname{\mathcal{O}}_{t}(A)=\int_{0}^{t}\mathds{1}_{A}(X_{s})d\langle X\rangle_{s},\qquad A=\text{Borel set of }\operatorname{\mathbb{R}}.$ (3)
See [50], [75, Chapter VI], and Section 2. The occupation flow $\operatorname{\mathcal{O}}=(\operatorname{\mathcal{O}}_{t})_{t\geq 0}$ is a measure-valued process adapted to the filtration of $X$. Moreover, the occupied process $(\operatorname{\mathcal{O}},X)$ is Markov provided that $X$ is Markov.
@@ 1.1 Contributions
$dX_{t}$ $=b(\operatorname{\mathcal{O}}_{t},X_{t})dt+\sigma(\operatorname{\mathcal{O}}_{t},X_{t})dW_{t},\hskip 17.00024ptd\operatorname{\mathcal{O}}_{t}=\delta_{X_{t}}\sigma(\operatorname{\mathcal{O}}_{t},X_{t})^{2}dt$ (6)
$dX_{t}$ $=b(\tilde{\operatorname{\mathcal{O}}}_{t},X_{t})dt+\sigma(\tilde{\operatorname{\mathcal{O}}}_{t},X_{t})dW_{t},\hskip 17.00024pt\ d\tilde{\operatorname{\mathcal{O}}}_{t}=\delta_{X_{t}}dt$ (7)
@@
$\Big(\partial_{\mathfrak{o}}+b\partial_{x}+\frac{\sigma^{2}}{2}\partial_{xx}\Big)u(\mathfrak{o},x)=0,\;\mathfrak{o}(\operatorname{\mathbb{R}})<T,\qquad u(\mathfrak{o},x)=\varphi(\mathfrak{o},x),\;\mathfrak{o}(\operatorname{\mathbb{R}})=T.$ (8)
@@
Proving the regularity of solutions of path-dependent PDEs is notoriously hard [44, 71] (see also [85, Chapter 11]) and still a subject of active research. In a recent work, Bouchard and Tan [17] study linear parabolic PPDEs for functionals depending on $X_{t}$ and $I_{t}=\int_{0}^{t}X_{s}dA_{s}$ for some continuous function $A$ of finite variation. Note that their setting relates to ours as the integral process $I$ can be retrieved from the occupation measure in some situations. Indeed, if $A_{t}\mathrel{\mathop{\ordinarycolon}}=\langle X\rangle_{t}$ is deterministic (e.g., when $X$ is a Gaussian martingale) and fulfills the conditions in [17], then $I_{t}=\int_{\operatorname{\mathbb{R}}}x\operatorname{\mathcal{O}}_{t}(dx)$.
@@
$\textnormal{v}(t,\omega)=\operatorname{\mathbb{E}}^{\operatorname{\mathbb{Q}}}[\varphi(\operatorname{\mathcal{O}}_{T},X_{T})|\operatorname{\mathcal{F}}_{t}](\omega),\qquad\omega\in\operatorname{\mathcal{D}}([0,T],\operatorname{\mathbb{R}}),$
@@ 1.2 Applications
Sections 4, 5, and 6 are dedicated to financial applications. We shall see that the occupation measure is ever-present, particularly in the eponymous occupation time derivatives [34, 62, 24], timer options, and corridor variance swaps. Other examples include Asian options contingent upon the running average, or Parisian options depending on the time spent by the asset above/below a barrier [24, 62]. It is a challenge for financial institutions to consistently price and manage complex derivatives books, and we show in 4 that occupation flows offer a unified framework for this task.
We then explore application of occupied processes in volatility modeling. In Section 5, we introduce the local occupied volatility (LOV) model that lies conveniently between local volatility [38] and path-dependent volatility [52, 53, 56, 59]. By design, this model guarantees a perfect fit to market prices of European vanilla options while providing additional flexibility to capture stylized facts of volatility or calibrate to other instruments such as VIX, barrier, and American options. We also propose a generalization of forward variance models in Section 6, opening the doors to nonlinear variance instruments contingent on the floating leg of corridor variance swaps.
@@ 2 Occupied Processes
$\operatorname{\mathcal{O}}_{t}(\omega)\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}})\to\operatorname{\mathbb{R}},\qquad\operatorname{\mathcal{O}}_{t}(\omega)(A)=\int_{0}^{t}\mathds{1}_{A}(X_{s})d\langle X\rangle_{s}(\omega),\qquad A\in\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}}),\qquad\omega\in\Omega.$ (9)
@@
    $\operatorname{\mathcal{O}}_{\tau}(A)=\operatorname{\mathcal{O}}_{\varsigma}(A)+\operatorname{\mathcal{O}}_{\varsigma,\tau}(A),\qquad\operatorname{\mathbb{Q}}-a.s.,$ (10)
@@
Evidently, the (strong) time additivity of $\operatorname{\mathcal{O}}$ carries over to $L$, that is $L_{\tau}=L_{\varsigma}+L_{\varsigma,\tau}$ with $L_{\varsigma,\tau}\mathrel{\mathop{\ordinarycolon}}=\frac{d\operatorname{\mathcal{O}}_{\varsigma,\tau}}{d\lambda}$ for all stopping times $\varsigma\leq\tau$. The local time is in fact central in the theory of additive functionals [14], [75, Chapter X]. For local martingales, it is well-known that the process $L_{t}=(L_{t}^{x})_{x\in\operatorname{\mathbb{R}}}$ (indexed by $x$) admits a version that is $\alpha$-Hölder continuous for all $\alpha<1/2$. For general semimartingales, $x\mapsto L_{t}^{x}$ may only shown to be càdlàg [75, Chapter VI]. Finally, note that $\operatorname{\mathcal{O}}_{t}(\operatorname{\mathbb{R}})=\langle X\rangle_{t}$, i.e., the total mass of $\operatorname{\mathcal{O}}$ coincides with the stochastic clock of $X$.
@@ 2.1 Calendar Time Occupation Flow
$\tilde{\operatorname{\mathcal{O}}}_{t}(A)\mathrel{\mathop{\ordinarycolon}}=\int_{0}^{t}\mathds{1}_{A}(X_{s})ds,\qquad A\in\operatorname{\mathcal{B}}(\operatorname{\mathbb{R}}),\qquad t\in[0,T],$ (14)
@@ 2.3 Occupation Functional and Markov Property
$W_{\theta}\mathrel{\mathop{\ordinarycolon}}=X_{\tau_{\theta}},\hskip 17.00024pt\tau_{\theta}=\inf\{t\in[0,T]\ \mathrel{\mathop{\ordinarycolon}}\ \langle X\rangle_{t}>\theta\},\hskip 17.00024pt\theta\leq\Theta\mathrel{\mathop{\ordinarycolon}}=\langle X\rangle_{T},$
@@
Note that $\theta_{\tau}$ is a stopping time under the Brownian filtration $\operatorname{\mathbb{F}}^{W}=(\operatorname{\mathcal{F}}_{\tau_{\theta}})_{\theta\leq\Theta}$. This observation will be particularly useful in Section 7 to derive simple formulation of optimal stopping problems. See also [72, Chapter IV].
@@ 2.4 Metrics on Space of Measures
$\operatorname{\mathfrak{m}}_{p}(\mathfrak{o},\mathfrak{o}^{\prime})\mathrel{\mathop{\ordinarycolon}}=\lVert\mathfrak{o}-\mathfrak{o}^{\prime}\rVert_{p},\qquad\lVert\mathfrak{o}\rVert_{p}\mathrel{\mathop{\ordinarycolon}}=\left\lVert\frac{d\mathfrak{o}}{d\lambda}\right\rVert_{L^{p}(\operatorname{\mathbb{R}})},\qquad\mathfrak{o},\mathfrak{o}^{\prime}\in\operatorname{\mathcal{M}}^{\lambda},\qquad p\in[1,\infty].$ (18)
@@
$\operatorname{\mathfrak{m}}_{p}(\operatorname{\mathcal{O}}_{s},\operatorname{\mathcal{O}}_{t})=\lVert\operatorname{\mathcal{O}}_{s,t}\rVert_{p}=\lVert L_{s,t}\rVert_{L^{p}(\operatorname{\mathbb{R}})},\qquad s<t.$
@@ 3.1 Occupation Derivative
$df(\operatorname{\mathcal{O}}_{t})=\delta_{\mathfrak{o}}f(\operatorname{\mathcal{O}}_{t})\cdot d\operatorname{\mathcal{O}}_{t},\qquad f\in\operatorname{\mathcal{C}}^{1}(\operatorname{\mathcal{M}}),$ (19)
@@
$\phi\cdot\mathfrak{o}\mathrel{\mathop{\ordinarycolon}}=\int_{\operatorname{\mathbb{R}}}\phi(x)\mathfrak{o}(dx),\qquad\phi\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}},\qquad\mathfrak{o}\in\operatorname{\mathcal{M}}.$ (20)
@@
$\delta_{\mathfrak{o}}(\mathfrak{o}\mapsto\phi\cdot\mathfrak{o})\equiv\phi,\qquad\forall\phi\in\operatorname{\mathcal{C}}(\operatorname{\mathbb{R}}).$ (22)
@@
$\partial_{\mathfrak{o}}f(\mathfrak{o},x)=\begin{cases}\delta_{\mathfrak{o}}f(\mathfrak{o})(x),\qquad&f\in\operatorname{\mathcal{C}}^{1}(\operatorname{\mathcal{M}}),\\[4.25006pt] \delta_{\mathfrak{o}}f(\mathfrak{o},x)(x),\qquad&f\in\operatorname{\mathcal{C}}^{1,0}(\operatorname{\mathcal{M}}\times\operatorname{\mathbb{R}}),\end{cases}$ (24)
@@ 3.2 Itô’s Formula
where $\nabla,\nabla^{2}$ denote the gradient and Hessian operator. The occupation derivative is again given by $\partial_{\mathfrak{o}}f(\mathfrak{o},x)=\delta_{\mathfrak{o}}f(\mathfrak{o},x)(x)$. A general Itô’s formula with arbitrary stochastic clock can be found in [78, Proposition 3.3].
@@ 3.2.1 Connection with the Functional Itô Calculus
$d\textnormal{f}(t,\omega)=\Delta_{t}\textnormal{f}(t,\omega)dt+\Delta_{x}\textnormal{f}(t,\omega)dX_{t}(\omega)+\frac{1}{2}\Delta_{xx}\textnormal{f}(t,\omega)d\langle X\rangle_{t}(\omega),\qquad\operatorname{\mathbb{Q}}-a.s.,$ (32)
@@ 3.2.2 Connection with the Malliavin Calculus
$D_{t}\textnormal{f}(\omega)\mathrel{\mathop{\ordinarycolon}}=\frac{d}{dh}\textnormal{f}(\omega+h\mathds{1}_{[t,T]})\Big|_{h=0}=\int_{t}^{T}\partial_{x\mathfrak{o}}f(\operatorname{\mathcal{O}}_{T}(\omega),\omega_{s})ds,\qquad\operatorname{\mathbb{Q}}-a.s.$ (34)